A uniform horizontal beam of length 8 m and weight
Answer #1

here,

length , l = 8 m

weight , w1 = 200 N

theta = 53 degree

Tmax = 1200 N

let the mass of the person be m

taking moment of force about the pin connection

Tmax * sin(theta) * l - m * g * l - w1 * l/2 = 0

1200 * sin(53) - m * 9.8 - 200/2 = 0

m = 87.59 kg

the mass of the person is 87.59 kg

a.

when x = 0.75 m

let the tension in the cable be T

taking moment of force about the pin connection

T * sin(theta) * l - m * g * x - w1 * l/2 = 0

T * sin(53) * 8 - 87.59 * 9.8 * 0.75 - 200 * 4/2 = 0

T = 163.37 N

the tension in the cable is 163.37 N

b.

let the force exerted on beam by wall be f

fx - T * cos(theta) = 0

fx = 163.37 * cos(53)

fx = 98.32 N

fy - w1 - m*g + T*sin(theta) = 0

fy = 200 + 87.59 * 9.8 - 163.37 * sin(53)

fy = 927.9 N

f = sqrt(fx^2 + fy^2)

f = 933.1 N

the force exerted on the beam by the wall is 933.1 N